# supplementary

### v1

@@ 1 Introduction
Despite achieving performance comparable to human experts on many tasks, the deployment of deep neural networks still faces many challenges, especially in sensitive domains such as medical imaging [25]. One major challenge is Shortcut Learning; a phenomenon where models exploit the presence of spurious features that coincidentally correlate with labels in the training data, despite not being relevant to the underlying relationship of interest. For instance, a review of hundreds of models for diagnosing COVID-19 from chest radiographs discovered that reported accuracies were often inflated by shortcut learning, with age being the spurious feature, and that none of the models can generalize to real world diagnostic settings [26]. Both the challenges of detecting when the model is learning shortcuts that rely on spuriously correlated (as opposed to core) features in the dataset, and of mitigating this shortcut learning, are active research topics that are also essential to the more general areas of fairness and bias mitigation in machine learning [3]. Despite recent progress, shortcut learning research still faces a number of challenges: First, Shortcut mitigation approaches often require knowledge of which features are spurious. In addition, many methods require group annotations for a subset of the data, i.e. knowledge of both the class labels and the presence / absence of the spurious feature [20, 24, 16]. However, this presupposition rarely holds in real-world settings. Secondly, methods that do not require group annotations often suppress various features without user input [33, 29]. However, use in sensitive domains requires the development of methods that allow users to retain control over the behavior of the model. Thirdly, many current methods involve modifying the data and retraining the model [20, 33, 24, 2]. These processes are highly computationally expensive. However, for a shortcut mitigation method to be widely adopted it should support easy exploratory analysis of model behavior when suppressing different correlations.
In this paper, we introduce a framework for effective, efficient and explainable shortcut detection and mitigation in transformers. Our approach leverages recent advancements in explainable AI, including prototype learning [6, 23] and Multi-Modal Large Language Models (MLLMs) concept identification [31], to enable robust shortcut detection and mitigation with minimal human supervision.
@@
The framework employs a multi-step process that analyzes model activations, detects prototypes in image patches, and uses MLLMs for prototype interpretation. The resulting interpretable shortcut detection component is then used to mitigate shortcuts during inference by selectively ablating image patches containing spurious features. Overall, our approach yields significant improvements in both worst group and average accuracy, offering a promising new direction in addressing shortcut learning. Furthermore, our framework can be extended to integrate with other shortcut mitigation techniques, such as Deep Feature Reweighting (DFR) [16]. The resulting symbiosis remedies several of DFR’s limitations [18] while improving performance. Finally, we conduct a user study to confirm that the discovered prototype concepts provide meaningful insight into the spurious features present in the data. In summary, our contributions are the following:
@@
  • We conduct a user study to demonstrate that the shortcut detection component of the framework learns human-friendly prototype concepts that enable users to correctly identify dataset shortcuts.
  • We show that the framework achieves competitive performance with state-of-the-art techniques that require group annotation. Furthermore, the framework can be easily integrated with other shortcut mitigation techniques, alleviating several of their inherent weaknesses, and achieving new state-of-the-art performance.
To encourage further experimentation, all code and data necessary to replicate the experiments discussed in this paper is made publicly available111github.com/Arsu-Lab/Shortcut-Detection-Mitigation-Transformers.
@@ 2 Related Work
Despite recent advancements in machine learning, including the rise of transformer-based architectures that surpass convolutional neural networks as the state-of-the-art, shortcut learning remains a persistent challenge [15]. Many studies tackling shortcut learning assume an ideal scenario where spurious features are known and group annotations are available [20, 24, 16]. However, recent evidence suggests that even widely-used datasets, such as ImageNet, contain previously unidentified shortcuts [19]. With this in mind, we explicitly tackled both shortcut detection and mitigation, without presupposing any prior knowledge.
@@ 2.1 Shortcut Detection
Many approaches to shortcut detection focus on gradient-based methods, such as Grad-CAM [27], to identify spurious features. For instance, DeGrave et al. [12] used gradients to create saliency maps, highlighting pixel importance across different COVID-19 X-ray images, and Zech et al. [32] applied activation maps to identify areas in chest radiographs that strongly influence model decisions. Although these methods have shown some success, they have been criticized for failing to provide the necessary insights to help users reliably identify spurious features [1].
One alternative approach suggested by Mueller et al[22] uses Variational Autoencoders to detect latent space dimensions with high label predictiveness and generate a set of images that differ in a single image attribute. The user then manually inspect these image sets to detect the attribute being manipulated and make a decision regarding it being a core or spurious feature.
@@ 2.2 Shortcut Mitigation
Under the ideal scenario, in which the spurious features are known and group annotation (shortcut presence and label) is available for at least a subset of the data, it is possible to edit model representations to eliminate reliance on encoded spurious attributes. For instance, Kirichenko et al. [16] proposed Deep Feature Reweighting (DFR), which uses explainability methods to identify and deactivate neurons in the last layer of a network that encode the spuriously correlated attribute. However, follow up work demonstrated that this may not be sufficient, as most neurons encode a combination of core and spurious attributes [18], suggesting that more exhaustive techniques are necessary to crowbar spurious and core attributes apart. For instance Liu et al. [20] uses extensive retraining while up-weighing samples from the worst performing group. While these methods achieve very high performance, their reliance on group annotations hinders their deployment in real world settings.
More recently, various approaches that alleviate this reliance on group annotations were developed [33, 2, 29]. Asgari et al. used saliency maps to identify which parts of the image the classifier depended on to reach its decisions, and masked these parts in future training to encourage model exploration, thereby mitigating model reliance on any one specific attribute [2]. Tiwari et al. [29] rely on a simple observation: spurious attributes are easier to learn relative to core attributes, and are learned by early model layers [29]. Therefore, by applying a loss to early layers that penalizes high classification accuracy the model is encouraged to forget the simple spurious attributes. Similarly, Zhang et al. [33] used the image labels to eliminate representation-space variability between samples belonging to the same class that resulted in sample misclassification. While these methods improve performance, they do not provide insights towards understanding what the model relies on. Moreover, considering the task-dependent nature of the status of an attribute as spurious or core, blind feature suppression might be detrimental to overall task performance.
@@ 2.3 Interpretable AI
Similarly to  [22], our approach also utilizes clustering in the representation space to detect images containing shortcuts. However, instead of using generative vision models, our work builds on ‘Prototype learning‘ approaches such as [6, 23]. These methods focus on learning human interpretable concept representations that allow classifiers to make decisions that can be justified to the end-user. In this vein, Yang et al. [31] proposed an innovative method that leverages MLLMs to generate semantic prototype descriptions and detect ‘bottleneck concepts‘ that uniquely identify specific labels.
Figure 1: An overview of our approach. A) a vision transformer is fine-tuned to perform the classification task. B) the representations learned by the last block are used to cluster the images, key-space embeddings for each image patch (blue rectangles) are saved to be used in later steps C) the most prototypical (closest to the cluster centroid) images are extracted. D) key-space distances to other patches in the same and other clusters are used to identify prototypical patches for every cluster. E) multi-modal LLMs generate descriptions of prototypical patches in each cluster which are further distilled into unique cluster concepts. F) domain expert in the loop uses these concepts to identify clusters containing spuriously correlated attributes. Alternatively, within cluster label homogeneity, within/between cluster distance of the prototypical patches, and Brier scores can be used for a completely unsupervised framework. G) in inference time, similarity to the key-space embeddings of the prototypical patches are used to identify if spurious attributes are contained in test image patches. H) Patches suspected of containing these attributes are then ablated.
@@ 3 Approach
For our baseline we used a ViT B-16 network pretrained on ImageNet-1K, that was fine-tuned on the training split of each of the datasets. We then execute our shortcut detection and mitigation pipeline using a holdout validation set. The first step of our approach is the extraction of patch activations from the final transformer block. These representations then undergo dimensionality reduction and clustering to identify salient patterns in the data. Within each cluster, we select representative samples based on their proximity to the cluster centroid, providing a representative subsample of the cluster’s composition.
A key innovation in our approach is the identification of prototypical patches within these representative samples. We achieve this through a meticulous examination of token distances in the key space. This approach allows us to pinpoint the most influential regions within each image.
We then employ a state-of-the-art multimodal large language model (MLLM) to generate captions for these prototypical patches and comprehensive summaries of the patches within each cluster. While these summaries can facilitate expert human review, we also introduce a simple yet effective heuristic for detecting shortcut clusters. This heuristic enables the potential for unsupervised operation of our pipeline, reducing the need for human intervention. Our method includes an ablation strategy that targets image tokens similar to the prototypical patches during inference.
@@ 3.1 Clustering
where $h_{L}^{t}(x_{i})$ is the embedding of the $t$-th token in the last layer $L$ for image $x_{i}$, and $T$ is the number of tokens (excluding the $<CLS>$ token). We apply PCA to reduce dimensionality from $d$ to $k$ (where $k=50$ in our experiments). We then apply unsupervised K-means clustering to these reduced embeddings. In contrast to Sohoni et al. [28] we do not need to apply over-clustering and settle for the naive approach of clustering into two clusters. We verified the effectiveness of this approach against overclustering (with $K$ selected by the silhouette method) and found no benefits (see Appendix 4).
@@ 3.2 Prototypical Patch Identification
We focus on a representative subset of samples collected from each cluster by taking the $N$ nearest samples to the centroid (we report results for $N=20$). Given these samples, our next step is to identify prototypical patches. These image patches should capture elements that are shared between images within each cluster, but are unique across different clusters (see Figure 1 C, D, and E).
@@ 3.2.1 Patch Similarity
Following Bolya et al. [4], instead of using image and intermediate model features, which can be over-parameterized, we employ the keys learned in the last self-attention block of the ViT model for each patch. This is, we mean over the keys of the 12-heads to reduce dimensionality, obtaining a summary of the information in each image token. Let $K(p)$ be the mean of the attention heads’ keys for patch $p$. We then compute the mean Euclidean distance $D(p)$ between each token and all tokens from other clusters:
$D(p)=\frac{1}{|C_{other}|}\sum_{q\in C_{other}}||K(p)-K(q)||_{2}$ (2)
where $C_{other}$ is the set of patches from all other clusters. To identify prototypical-patches that are both unique for each cluster and important for the model’s decision making, we compute a prototypicality score $P(p)$ for each patch. Patches with high $P(p)$ scores are considered prototypical for their respective clusters (figure 1 E). We extract the top $M$ patches from the selected subset of each cluster for the following steps (we report results for $M=200$).
@@ 3.3 Spurious Concept Identification
Given the patches we identified as prototypical in the previous step we extract them from the pixel space and use them to generate human understandable concepts. This is achieved using pretrained publicly available MLLMs. Specifically, we first use LLaVA [21], a model that bridges language and vision by learning a projection from the pretrained visual encoder of CLIP to Vicuna, an open source large language model [9]. We use LLaVA to extract text description from prototypical patches for each cluster. The output is further refined to find the semantic contexts that best capture each cluster. This can be achieved efficiently by using KL divergence to compute concept bottlenecks similarly to previous work [31]. We found that instructing a large language model to complete this refinement step generated superior performance. We tested three different LLMs and found excellent performance even when using the smallest LLaMA model [30]. The text output enabled users to easily identify which cluster contains a shortcut (see user study in section 5).
@@ 3.3.1 Unsupervised Spurious Concept Identification
Alternative to the prior step we developed a simple heuristic to identify a shortcut cluster, which we use in all our experiments. While we believe human feedback is crucial for correct model behavior, this unsupervised approach could be useful when conducting cursory exploratory analysis. For unsupervised cluster selection we used a combination of the homogeneity and Brier score. The cluster homogeneity score $h_{c}$ measures how much the cluster contains samples from a single class:
@@
where $f_{i}$ is the predicted probability, $o_{i}$ is the actual outcome (0 or 1), and $N$ is the number of predictions. The intuition behind this approach is the following: spuriously correlated attributes are, by definition, vastly over represented in a single class. Therefore, we expect the cluster representing the spuriously correlated attributes to have high homogeneity. Similarly, the existence of the spurious feature is a strong signal of class membership. Therefore, as observed in previous research, the classifier will make confident predictions (high Brier score) for images containing the spurious attribute [29]. Since these predictions are only correct for the dominant class we calculate two Brier scores: one for the dominant ($bd_{c}$) and one for the non-dominant ($bn_{c}$) class samples. We expect highly confident correct predictions for the dominant class samples leading to a low Brier Score and highly confident incorrect predictions for the non-dominant class samples (since it contains the spurious attribute but is not part of the spurious correlated class). Hence, given the average homogeneity score $h_{c}$ and dominant Brier score $bd_{c}$ as well as non-dominant Brier score $bn_{c}$ per cluster, our unsupervised method picks the cluster that maximizes a combination of the three:
$\underset{c\in C}{\mathrm{argmax}},{\lambda_{1}h_{c}+\lambda_{2}e^{-bd_{c}}+\lambda_{3}(1-e^{-bn_{c}})}$
Where $\sum\lambda_{i}=1$ (we weigh all factors equally in the experiments). We use the exponent of the Brier scores to map all equation elements to the same range. As we discuss in section 5.3, a user study found that judgements of participants overwhelmingly aligned with the predictions of this unsupervised method.

### v2

@@ 1 Introduction
Despite achieving performance comparable to that of human experts in many tasks, the deployment of deep neural networks still faces many challenges, especially in sensitive domains such as medical imaging [26]. One major challenge is Shortcut Learning; a phenomenon where models exploit the presence of spurious features that coincidentally correlate with labels in the training data, despite not being relevant to the underlying relationship of interest. For instance, a review of hundreds of models for diagnosing COVID-19 from chest radiographs discovered that reported accuracies were often inflated by shortcut learning, with age being the spurious feature, and that none of the models can generalize to real world diagnostic settings [27]. Both the challenges of detecting when the model is learning shortcuts that rely on spuriously correlated (as opposed to core) features in the dataset, and of mitigating this shortcut learning, are active research topics that are also essential to the more general areas of fairness and bias mitigation in machine learning [3]. Despite recent progress, shortcut learning research still faces a number of challenges: First, Shortcut mitigation approaches often require knowledge of which features are spurious. In addition, many methods require group annotations for a subset of the data, i.e., knowledge of both the class labels and the presence / absence of the spurious feature [21, 25, 17]. However, this assumption rarely holds in real-world settings. Secondly, in methods that do not require group annotations, the user often has no knowledge of the features were suppressed [34, 30]. However, use in sensitive domains requires the development of methods that allow users to retain control over the behavior of the model. Third, many current methods involve modifying the data and retraining the model [21, 34, 25, 2]. These processes are highly computationally expensive. However, for a shortcut mitigation method to be widely adopted, it should support easy exploratory analysis of model behavior when suppressing different correlations.
@@
In this paper, we introduce a framework for effective, efficient, and explainable shortcut detection and mitigation in transformers. Our approach leverages recent advancements in explainable AI, including prototype learning [6, 24] and Multi-Modal Large Language Models (MLLMs) concept identification [32], to enable explainable unsupervised shortcut detection and mitigation. The framework employs a multi-step process that analyzes model activations, detects prototypes in image patches, and uses MLLMs for prototype interpretation. The resulting interpretable shortcut detection component is then used to mitigate shortcuts during inference by selectively ablating image patches containing spurious features. Overall, our approach yields significant improvements in both worst group and average accuracy, offering a promising new direction in addressing shortcut learning. Finally, we conduct a user study to confirm that the prototype concepts discovered provide a meaningful insight into the spurious features present in the data. In summary, our contributions are the following:
@@
  • We show that the framework outperforms state-of-the-art techniques that require group annotation. Furthermore, the framework eliminates shortcuts from the stimuli directly, alleviating concerns regarding any remanence of shortcut learning [19].
  • We conduct a user study to demonstrate that the shortcut detection component of the framework learns human-friendly prototype concepts that enable users to correctly identify dataset shortcuts, making our framework interpretable, in contrast to previous unsupervised methods.
To encourage further experimentation, all code and data necessary to replicate the experiments discussed in this paper are made publicly available111github.com/Arsu-Lab/Shortcut-Detection-Mitigation-Transformers.
@@ 2 Related Work
Despite recent advances in machine learning, including the rise of transformer-based architectures that surpass convolutional neural networks as the state-of-the-art, shortcut learning remains a persistent challenge [15]. Many studies tackling shortcut learning assume an ideal scenario in which spurious features are known and group annotations are available [21, 25, 17]. However, recent evidence suggests that even widely used datasets, such as ImageNet, contain previously unidentified shortcuts [20]. With this in mind, we explicitly tackled both shortcut detection and mitigation, without presuming any prior knowledge.
@@ 2.1 Shortcut Detection
Many approaches to shortcut detection focus on gradient-based methods, such as Grad-CAM [28], to identify spurious features. For example, DeGrave et al. [12] used gradients to create saliency maps, highlighting pixel importance across different COVID-19 X-ray images, and Zech et al. [33] applied activation maps to identify areas in chest radiographs that strongly influence model decisions. Although these methods have shown some success, they have been criticized for failing to provide the necessary insights to help users reliably identify spurious features [1].
An alternative approach suggested by Mueller et al[23] uses Variational Autoencoders to detect latent space dimensions with high label predictiveness and generate a set of images that differ in a single image attribute. The user then manually inspects these image sets to detect the attribute being manipulated and makes a decision as to whether it is a core or spurious feature.
@@ 2.2 Shortcut Mitigation
Under the ideal scenario, in which the spurious features are known and group annotation (shortcut presence and label) is available for at least a subset of the data, it is possible to edit model representations to eliminate reliance on encoded spurious attributes. For instance, Kirichenko et al. [17] proposed Deep Feature Reweighting (DFR), which utilizes last layer retraining with a group-wise balanced dataset to effectively deactivate neurons that encode the spuriously correlated attribute. However, follow up work demonstrated that this may not be sufficient, as most neurons encode a combination of core and spurious attributes [19], suggesting that more exhaustive techniques are necessary to crowbar spurious and core attributes apart. For instance, Liu et al. [21] uses extensive retraining while up-weighing samples from the worst performing group. Similarly, Idrissi et al. [16] proposed SUBG (subsampling groups), which reduces all groups to the size of the smallest group by randomly subsampling examples from larger groups, effectively removing group imbalance during training. While these methods achieve very high performance, their reliance on group annotations hinders their deployment in real-world settings.
More recently, various approaches were developed that alleviate this dependency on group annotations [34, 2, 30]. Asgari et al. used saliency maps to identify which parts of the image the classifier depended on to reach its decisions and masked these parts in future training to encourage model exploration, thereby mitigating model reliance on any one specific attribute [2]. Tiwari et al. [30] rely on a simple observation: spurious attributes are easier to learn relative to core attributes and are learned by early model layers [30]. Therefore, by applying a loss to early layers that penalizes high classification accuracy, the model is encouraged to forget the simple spurious attributes. Similarly, Zhang et al. [34] used the image labels to eliminate representation-space variability between samples belonging to the same class that resulted in sample misclassification. While these methods improve performance, they do not provide insights towards understanding what the model relies on. Moreover, considering the task-dependent nature of the status of an attribute as spurious or core, blind feature suppression might be detrimental to overall task performance.
@@ 2.3 Interpretable AI
Similarly to  [23], our approach also utilizes clustering in the representation space to detect images containing shortcuts. However, instead of using generative vision models, our work builds on ‘Prototype learning‘ approaches such as [6, 24]. These methods focus on learning human interpretable concept representations that allow classifiers to make decisions that can be justified to the end user. In this vein, Yang et al. [32] proposed an innovative method that leverages MLLMs to generate semantic prototype descriptions and detect ‘bottleneck concepts‘ that uniquely identify specific labels.
Figure 1: An overview of our approach. A) a vision transformer is fine-tuned to perform the classification task. B) the representations learned by the last block are used to cluster the images, key-space embeddings for each image patch (blue rectangles of the ViT) are saved to be used in step G. C) the most prototypical (closest to the cluster centroid) images are extracted. D) key-space distances to other patches in the same and other clusters are used to identify prototypical patches for every cluster. E) multi-modal LLMs generate descriptions of prototypical patches in each cluster which are further distilled into unique cluster concepts. F) domain expert in the loop uses these concepts to identify clusters containing spuriously correlated attributes. Alternatively, cluster label homogeneity, within/between cluster distance of the prototypical patches, and Brier scores can be used for a completely unsupervised framework. G) Similarity to the key-space embeddings of the prototypical patches are used to identify if spurious attributes are contained in validation image patches. H) Patches with high similarity to prototypical spurious patches are ablated, and the last linear layer is retrained to make decisions based on core features.
@@ 3 Approach
For our baseline, we used a ViT B-16 network pre-trained on ImageNet-1K that was fine-tuned on the training split of each of the datasets. We then execute our shortcut detection and mitigation pipeline using a holdout validation set. The first step of our approach is to extract the patch activations from the final transformer block. These representations then undergo dimensionality reduction and clustering to identify salient patterns in the data. Within each cluster, we select representative samples based on their proximity to the cluster centroid, providing a representative subsample of the cluster’s composition.
A key innovation in our approach is the identification of prototypical patches within these representative samples. We achieve this through a meticulous examination of token distances in the key space. This approach allows us to pinpoint the most influential patches within each image. We then employ a multimodal large language model (MLLM) to generate captions for these prototypical patches and comprehensive summaries of the patches within each cluster. These summaries allow a human expert to review - and if necessary intervene in - decisions made by the unsupervised pipeline, which employs simple yet effective heuristics to detect shortcut clusters. Our method then applies an ablation strategy that targets image tokens similar to those in the prototypical patch cluster identified as containing shortcuts. Finally, the last model layer is retrained to reinforce the reliance on core features during classification.
As demonstrated in Figure 1, the proposed approach is highly modular, consisting of a series of independent components. Although multiple alternatives for each component achieved comparable results, this set-up was ultimately chosen due to computational efficiency and simplicity.
@@ 3.1 Clustering
where $h_{L}^{t}(x_{i})$ is the embedding of the $t$-th token in the last layer $L$ for image $x_{i}$, and $T$ is the number of tokens (excluding the $<CLS>$ token). We apply PCA to reduce the dimensionality from $d$ to $k$ (where $k=50$ in our experiments). We then apply unsupervised K-means clustering to these reduced embeddings. In contrast to Sohoni et al. [29], we do not apply overclustering, instead opting for the naive approach of clustering into two clusters. We verified the effectiveness of this approach against overclustering (with K selected by the silhouette method) and found no benefits, while also confirming clustering accuracy (see Appendix D).
@@ 3.2 Prototypical Patch Identification
We focus on a representative subset of samples collected from each cluster by taking the $N$ nearest samples to the centroid (we report results for $N=20$). Given these samples, our next step is to identify prototypical patches. These image patches should capture elements that are shared between images within each cluster, but are unique between different clusters (see Figure 1 C, D, and E).
@@ 3.2.1 Patch Similarity
Following Bolya et al. [4], instead of using the intermediate and image model features, which can be overparameterized, we employ the keys learned in the last self-attention block of the ViT model for each patch. This is, we mean over the keys of the 12-heads to reduce dimensionality, obtaining a summary of the information in each image token. Let $K(p)$ be the mean of the attention heads’ keys for patch $p$. To identify prototypical-patches that are both unique for each cluster and important for the model’s decision making, we compute a prototypicality score $P(p)$ for each patch, which is the Euclidean distance between each token and all tokens from other clusters:
$P(p)=\frac{1}{|C_{other}|}\sum_{q\in C_{other}}||K(p)-K(q)||_{2}$ (2)
where $C_{other}$ is the set of patches from all other clusters. Patches with high $P(p)$ scores are considered prototypical of their respective clusters (Figure 1 E). We extract the top $M$ patches from the selected subset of each cluster for the following steps (we report results for $M=200$).
@@ 3.3 Spurious Concept Identification
Given the patches we identified as prototypical in the previous step, we extract them from the pixel space and use them to generate human-understandable concepts. This is achieved using publicly available pre-trained MLLMs. Specifically, we first use LLaVA [22], a model that bridges language and vision by learning a projection from the pretrained visual encoder of CLIP to Vicuna, an open source large language model [9]. We use LLaVA to extract text description from prototypical patches for each cluster. The output is further refined to find the semantic contexts that best capture each cluster. This can be achieved efficiently using KL divergence to compute concept bottlenecks similar to previous work [32]. However, we found that instructing a large language model to complete this refinement step generated superior performance. We tested three different LLMs and found excellent performance even when using the smallest LLaMA model [31]. The text output allowed users to easily identify which cluster contains a shortcut (see the user study in Section 5). While human input is not necessary, and this step can be skipped, the interpretability afforded by this step is crucial for real world applications.
@@ 3.3.1 Unsupervised Spurious Concept Identification
Alternative to the prior step we developed a simple heuristic to identify a shortcut cluster, which we use in all our experiments. Although we believe human feedback is crucial for correct model behavior, this unsupervised approach could be useful when conducting cursory exploratory analysis. For unsupervised cluster selection, we used a combination of homogeneity and Brier score. The cluster homogeneity score $h_{c}$ measures how much the cluster contains samples from a single class:
@@
where $f_{i}$ is the predicted probability, $o_{i}$ is the actual outcome (0 or 1), and $N$ is the number of predictions. The intuition behind this approach is the following: spuriously correlated attributes are, by definition, vastly overrepresented in a single class. Therefore, we expect that the cluster representing the spuriously correlated attributes will have high homogeneity. Similarly, the existence of the spurious feature is a strong signal of class membership. Therefore, as observed in previous research, the classifier will make confident predictions for images containing the spurious attribute [30]. Since these predictions are only correct for the dominant class, we calculate two Brier scores: one for the dominant ($bd_{c}$) and one for the nondominant ($bn_{c}$) class samples. We expect highly confident correct predictions for the dominant class samples leading to a low Brier score and highly confident incorrect predictions for the nondominant class samples (since it contains the spurious attribute but is not part of the spurious correlated class) leading to a high Brier score. Hence, given the average homogeneity score $h_{c}$ and dominant Brier score $bd_{c}$ as well as non-dominant Brier score $bn_{c}$ per cluster, our unsupervised method picks the cluster that maximizes a combination of the three:
$\underset{c\in C}{\mathrm{argmax}},{\lambda_{1}h_{c}+\lambda_{2}e^{-bd_{c}}+\lambda_{3}(1-e^{-bn_{c}})}$ (5)
Where the exponent of the Brier scores is used to map all elements to the same range. In our experiments, we weigh the three elements equally. As we discuss in Section 5.3, a user study found that human judgments overwhelmingly aligned with the predictions of this unsupervised method.